Suppose Jones Company has orders from three customers located in the same market area. One order has a total weight of 4,300 pounds, the second weighs 9,000 pounds, and the third weighs 12,800 pounds. The transportation carrier quotes a freight rate of $18 per hundredweight (or cwt.) for direct shipment to the customer for shipments weighing 1,000 to 4,999 pounds, $16 per cwt. for orders weighing 5,000 to 9,999 pounds, and $14 for shipments weighing between 10,000 and 15,000 pounds. Alternatively, the carrier’s rate for shipments weighing more than 20,000 pounds is $10.00 per cwt. However, if the orders are combined into one consolidated shipment, the carrier will charge $330 for each stop it is required to make. a. Calculate the total cost using separate shipments, and total cost using consolidated shipments
Answer #1

Total Cost using Seperate Shipments

For Order 1- Cost Per hundredweight * Number of hundred weight = 18 * 4300 / 100 = 774

For Order 2- Cost Per hundredweight * Number of hundred weight = 16 * 9000 / 100 = 1440

For Order 3- Cost Per hundredweight * Number of hundred weight = 14 * 12800 / 100 = 1792

For Consolidated Shipment-

Number of Stops to be made = 2 (Only Two Intermediate Stops are considered)

Total Weight of Shipment = 4300+9000+12800 = 26100

Total Cost for consolidated order = Cost Per hundredweight * Number of hundred weight + Number of Stops * Cost per Stops = 10 * 26100 / 100 + 2 * 300 = 32100

Please Like and provide your reviews in comments. :-)